OpenStudy (anonymous):

how to set this up and solve for it? the ratio of men to women on a bus was 5/7. Then two women and one man boarded, and the ration was 7/10. How many men and women were on the bus before the last three passengers boarded?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@robtobey

OpenStudy (phi):

first step, label the number of men x and women y the ratio of men to women on a bus was 5/7 means \[ \frac{x}{y}= \frac{5}{7} \] Then two women and one man boarded Now how many men do we have? How many women do we have?

OpenStudy (anonymous):

1/2=5/7 so we would have 10 women and 7 men

OpenStudy (phi):

I am sure you know ½ does not equal 5/7 think like this: I start with x men. One man boards the bus. There is now x plus ? men on board ?

OpenStudy (anonymous):

i know what i did was multiply them to get 7/10

OpenStudy (phi):

do you see we have x+1 men on board ? now how many women ?

OpenStudy (anonymous):

y+2

OpenStudy (anonymous):

for women

OpenStudy (phi):

Then two women and one man boarded, and the ratio was 7/10. and what ratio can we set up ? (x+1)/(y+2) = ?

OpenStudy (phi):

you should get \[ \frac{x+1}{y+2}= \frac{7}{10} \] cross multiply to get \[ 10x +10= 7y+14 \\ 10x -7y= 4 \] you also have the original ratio \[ \frac{x}{y}= \frac{5}{7}\\ 7x= 5y\\ 7x-5y=0\] you have 2 equations and 2 unknowns \[ 7x-5y=0\\10x -7y= 4\] can you solve for x and y ?

OpenStudy (anonymous):

sorry i had to do something

OpenStudy (anonymous):

yes i can solve for this...

OpenStudy (anonymous):

i would get 20 for x amount of men and 28 y amount of women x=20,y=28

OpenStudy (phi):

and 20/28 simplifies (divide top and bottom by 4) to 5/7 which matches the question add 1 man and 2 women to get 21/30 which simplifies to 7/10. The numbers make sense.

OpenStudy (anonymous):

yes